OpenStudy (anonymous):

Simplify. √108 - 3√8 - √27

OpenStudy (anonymous):

√108-3√8-√27 = sqrt(36 * 3) - 3sqrt(4 * 2) - sqrt(9 * 3) = 6sqrt(3) - 3(2sqrt(2)) - 3sqrt(3) = 6sqrt(3) - 6sqrt(2) - 3sqrt(3) = 3sqrt(3) - 6sqrt(2) c) 3√3-6√2

OpenStudy (anonymous):

that's what i got am i correct? ^^^^^

OpenStudy (whpalmer4):

Yep, looks good!

OpenStudy (whpalmer4):

Here I was writing you an explanation, but you've already mastered it!

OpenStudy (anonymous):

thank you!